Boxell associates has $7,400 invested at 5.5% simple interest. After money was withdrawn, $242 was earned for one year on the remaining funds. How much was withdrawn?
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

7400 is invested at 5.5% simple interest.
x equals amount withdrawn.
(7400 - x) * .055 = 242.
simplify to get:
7400 * .055 - .055 * x = 242
simplify further to get:
407 - .055 * x= 242
subtract 242 from both sides of the equation and add .055 * x to both sides of the equation to get:
407 - 242 = .055 * x
combine like terms to get:
165 = .055 * x
solve for x to get:
x = 165 /.055 = 3000.
looks like 3000 was withdrawn.
7400 - 3000 = 4400
4400 * .055 = 242.
amount withdrawn is confirmed to be correct.
your solution is that 3000 was withdrawn.

Answer by ikleyn(52781)

.

Not a Math problem.


As it worded,  formulated and presented,  it is

            G I B B E R I S H


in its purest form.